Find a six-digit number in which the first digit is one less than the third, the third digit is one less
than the fifth, and the second digit is eight more than the first, one more than the fourth, and two more than the last. The sum of the first and fourth digits equals the second digit, the sum of the first and third digits equals the fifth, and the sum of all the digits is 30.

Answer by Edwin McCravy(20081)

Find a six-digit number
We put 6 blanks for the 6 digits:

_ _ _ _ _ _

the second digit is eight more than the first,
Since the first digit can't be 0, then in order for there to be a
digit which is 8 more than the first, the first digit is 1 and
the second digit is 9.  So we have:

1 9 _ _ _ _

(the second digit is) one more than the fourth,
So the fourth digit must be 8, so that 9 will be one more than it.

So we have:

1 9 _ 8 _ _

and (the second digit is) two more than the last.
So the last digit must be 7, so that 9 will be two more than it.

So we have:

1 9 _ 8 _ 7

in which the first digit is one less than the third,
So the third digit is 2.  So we have

1 9 2 8 _ 7

the third digit is one less than the fifth,
So the fifth digit is 3. So we have the entire number.

1 9 2 8 3 7

and The sum of the first and fourth digits equals the second digit,
That's true, 1+8=9, but we didn't need to know that.

the sum of the first and third digits equals the fifth,
That's true too, 1+2=3, but we didn't need to know that either.

and the sum of all the digits is 30.
That's also true, 1+9+2+8+3+7=30, but we didn't need to know that either.

Edwin